OpenStudy (winsteria):

Which fractions are greater than -7/12 ? Choose exactly two answers that are correct. A) -2/3 B) -2/5 C) -5/9 D) -13/18

OpenStudy (winsteria):

@TheSmartOne

OpenStudy (winsteria):

no sorry I'll go check

OpenStudy (winsteria):

This question has me confused :/

OpenStudy (winsteria):

Someone told me it was A and D, and someone else told me it was B and C?

OpenStudy (anonymous):

do you have a calculator

OpenStudy (winsteria):

well yeah

OpenStudy (anonymous):

ok so divide them all each

OpenStudy (anonymous):

like 2/3 . it will help you understand it more in decimals

OpenStudy (anonymous):

did you get them into decimals

OpenStudy (winsteria):

Yeah

OpenStudy (winsteria):

I've been on this question for like an hour idek

OpenStudy (anonymous):

ok can you type the out them out

TheSmartOne (thesmartone):

-7/12=-0.58333333333

TheSmartOne (thesmartone):

-2/3=-.6666 So this number is more negative so it is not the answer...

TheSmartOne (thesmartone):

-2/5=-.4 This is a smaller negative number so it is greater than -7/2 so it is one of the answer.

TheSmartOne (thesmartone):